OpenStudy (anonymous):

WILL MEADEL AND FAN!!!!!!!!!!!!!!!! Emily has 85 books on her book shelf. of those books, 17 are mysteries 18 are adventure stories, and the rest are fantasies what is the simplified ratio of fanitasies to total books on her bookshelf. A.35:50 B.7:17 C.5:8 D10:17

OpenStudy (anonymous):

What is 17+18?

OpenStudy (anonymous):

35

OpenStudy (anonymous):

Then after you find that out subtract it from 85 to get the number of fantasy books

OpenStudy (anonymous):

17 : 10

OpenStudy (anonymous):

Good, what is 85-35?

OpenStudy (anonymous):

PLease explain your answer :) Thanks!

OpenStudy (anonymous):

50

OpenStudy (anonymous):

Good so now we have 50/85 or \(\frac{50}{85}\)

OpenStudy (anonymous):

SO NOW WHAT DO I DO

OpenStudy (anonymous):

Let's find the highest common factor.. Factors of 50 - 1, 2, 5, 10, 25, 50 Factors of 85 - 1, 5, 17, 85 What is the highest number they have in common?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Great! Now we divide both numbers of the ratio by 5, so what is 50/5 and 85/5?

OpenStudy (anonymous):

50/5 : 85/5

OpenStudy (anonymous):

85 - 35 =50 85 : 50 ( divide both by 5) 17 :10

OpenStudy (anonymous):

<_>..

OpenStudy (anonymous):

ive been trying to post this for the longest and i coulnt post it

OpenStudy (anonymous):

It is 10:17

OpenStudy (anonymous):

your right :)

OpenStudy (anonymous):

Because 50/5 = 10 ad 85/5 is 17, so therefore we have 10:17